OpenStudy (anonymous):

What is the equation of the line, in slope-intercept form, that passes through (4, 2) and (-2, -3)? 5x - 6y - 8 = 0 3y = -2x + 20 y = 5/6x - 4/3

OpenStudy (anonymous):

m = y2-y1 / x2-x1 = -3-2 / -2-4 = -5/-6 = 5/6 (4,2) y2 - y1 = m(x2-x1) y - 2 = 5/6(x-4) y - 2 = 5/6x - 20/6 y = 5/6x -20/6 + 2 y = 5/6x -8/6 y = 5/6x -4/3 so last option C.

OpenStudy (anonymous):

ok thanks